OpenStudy (anonymous):

What is the volume of a sphere with a radius of 9 inches? 108π 242π 642π 972π

OpenStudy (kidrah69):

\[V=\frac{ 4 }{ 3 }\pi~r^3\]

OpenStudy (kidrah69):

\[V=\frac{ 4 }{ 3}(3.14)(9)^3\]

OpenStudy (kidrah69):

Solve.

OpenStudy (anonymous):

wait what

OpenStudy (kidrah69):

Ok first i showed the formula then all i did was plug in the numbers given.

OpenStudy (anonymous):

ok one min

OpenStudy (anonymous):

ok would it be c

OpenStudy (anonymous):

@kidrah69

OpenStudy (kidrah69):

No thats not right...

OpenStudy (anonymous):

do not plug he value of pi

OpenStudy (kidrah69):

Oh right >_< it doesnt even tell you the value of pi here! so just do whats in the formula and leave pi out

OpenStudy (anonymous):

A

OpenStudy (k_lynn):

That's not quite right either. The formula is \[\frac{ 4 }{ 3}\Pi r ^{3}\] However since there is \(\pi\) in your answer choices, do not multiply by pi in the equation. This will be the equation you will use. \[\frac{ 4 }{ 3 }r ^{3}\] The question gives you the radius. That's 9 inches. What the 4/3 means is divide. So When you break the formula down, this is the math you'll be doing. \[4\div3\times9^{3}\] \(9^3\) means to multiply 9 by itself 3 times. So like this. \[4\div3\times9\times9\times9\]Can you do that for me? :)

OpenStudy (anonymous):

947.7

OpenStudy (anonymous):

so D

OpenStudy (k_lynn):

Hmm. I think you messed up somewhere in your math. Do you have a scientific calculator?

OpenStudy (k_lynn):

Yep, D.

OpenStudy (k_lynn):

:)